# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

### v1

@@ 3.2 No-three-way interaction model of three-way table
is random, the vectors of the sufficient statistics is
@@ A Markov basis consists of [6]
First, we assessed the convergence of the generalized IPS. Unlike the case of the two-way contingency tables in Section 3.1, setting $\epsilon$ in equation (16) needed care. If the convergence was insufficient, a sample path hit outside the Markov lattice because of the numerical error in computing transition probabilities with the floating point arithmetic. The generalized IPS always converged if we took small $\epsilon$ but demanded many iterations. In the following numerical experiments, we set $\epsilon=0.005$ throughout. If we could not achieve convergence by 1,000 iterations (less than 0.06%), we discarded the sample path and drew another contingency table.
@@
and $p(z)=0$ if $z\notin\{0,8,12\}$. The closeness to the true distribution (22) of the chi-square values demonstrates that the MLE performs a good approximation of the UMVUE. The pairs $(10^{4},10^{4})$ and $(10^{5},10^{4})$ were equally good, which implies that the Metropolis chain converged after 100,000 steps. An important observation is that the total variation distance of the pair $(10^{4},10^{5})$ decreases as $s$ increase except $s=1$ (the Monte Carlo error was small for $s=1$ since we used not the empirical distribution $\hat{p}_{D}$ but the true distribution (22)). Unlike the case of the two-way contingency tables (Table 1), the bias decrease of the MLE is evident.
@@ 4 Discussion
Previous implementations of the direct sampling algorithm [14, 16] have enabled exact sampling, but depending on the model, the computational burden makes it difficult to put into practical use. In this study, we proposed the approximate algorithm applicable to any log-affine model with the expense of the bias, that asymptotically disappears. In contrast to deriving the Markov basis in the Metropolis algorithm or the connection matrix of the holonomic ideal in the exact sampling, we can implement the approximate sampling algorithm straightforwardly since it demands no prior investigations; it only needs existing methods to compute the MLE.
@@
There is room for improvement in the approximate direct sampling algorithm proposed in this paper. In terms of accuracy, it can approach UMVUE by reducing the bias of the MLE of the expected count, and in terms of the speed of computing the MLE, Newton’s method can be faster than the generalized iterative proportional scaling. However, these may be consideration when applying the algorithm to each model and are beyond the scope of this paper, which aims to propose an approximate sampling algorithm in its most general form.

### v2

@@ 3.2 No-three-way interaction model of three-way table
is random, the vector of the sufficient statistics is
@@ A Markov basis consists of [6]
First, we assessed the convergence of the generalized IPS. Unlike the case of the two-way contingency tables in Section 3.1, setting $\epsilon$ in equation (16) needed care. Proposition 2.13 guarantees that a sample path is on the Markov lattice, even for the approximate direct sampling algorithm. However, if the convergence was insufficient, a sample path hit outside the Markov lattice because of the numerical error in computing transition probabilities with the floating point arithmetic. We confirmed that the generalized IPS always converged if we took small $\epsilon$ but demanded many iterations. In the following numerical experiments, we set $\epsilon=0.005$ throughout. If we could not achieve convergence by 1,000 iterations (less than 0.06%), we discarded the sample path and drew another contingency table.
@@
and $p(z)=0$ if $z\notin\{0,8,12\}$. The closeness to the true distribution (22) of the chi-square values demonstrates that the MLE performs a good approximation of the UMVUE. The pairs $(10^{4},10^{4})$ and $(10^{5},10^{4})$ were equally good, which implies that the Metropolis chain converged after 100,000 steps. An important observation is that the total variation distance of the pair $(10^{5},10^{5})$ decreases as $s$ increase except $s=1$ (the Monte Carlo error was small for $s=1$ since we used not the empirical distribution $\hat{p}_{D}$ but the true distribution (22)). Unlike the case of the two-way contingency tables (Table 1), the bias decrease of the MLE is evident.
@@ 4 Discussion
Previous implementations of the direct sampling algorithm [14, 16] have enabled exact sampling, but depending on the model, the computational burden makes it challenging to put into practical use. In this study, we proposed the approximate algorithm applicable to any log-affine model with the expense of the bias, that asymptotically disappears. In contrast to deriving the Markov basis in the Metropolis algorithm or the connection matrix of the holonomic ideal in the exact sampling, we can implement the approximate sampling algorithm straightforwardly since it demands no prior investigations; it only needs existing methods to compute the MLE.
There is room for improvement in the approximate direct sampling algorithm proposed in this paper. In terms of accuracy, it can approach the UMVUE by reducing the bias of the MLE of the expected count, and in terms of the speed of computing the MLE, Newton’s method can be faster than the generalized iterative proportional scaling. However, these are typical tasks in statistical analysis when applying an algorithm to each model and are beyond the scope of this paper, which aims to propose an approximate sampling algorithm in its most general form.